# Tackling the Inherent Difficulty of Noise Filtering in RAG

### v1

@@ 1 Introduction
  • Filtering out irrelevant information is inherently difficult, limited number of transformer layers fail to solve it.
  • Fine-tuning the model can hardly distinguish irrelevant information while taking advantage of relevant ones.
About the difficulty of filtering out irrelevant information. Considering the query ‘Alice is exhausted but Bob is still excited; how does Bob feel?’ Here, ‘exhausted’ is noise and should be excluded from the inference. However, assessing relevance necessitates considering the token ‘Bob’ in the query alongside ‘Alice’ which is the subject of ‘exhausted’. Thus, evaluating relevance requires the information from three or even more tokens. Yet, the attention mechanism typically computes only pairwise relationships, making it challenging to resolve this issue within a limited number of transformer layers (3Match). Therefore filtering out noise documents is inherently difficult which explains why current retrieval methods (Contriever; BM25; DPR) would always incorporate some noisy information in the retrieval results.
Therefore, a question is the powerful LLMs much more parameters able to solve this. We hope that the LLM could be robust to noisy information while extracting helpful information (RAG_robust; RAG_robust_finetune_25). However, we argue that standard fine-tuning is structurally ill-suited for this task. The core issue lies in a fundamental trade-off imposed by their linear update mechanism. To filter noise, the learned parameter update must apply a strong negative adjustment to the attention scores of irrelevant tokens. Yet, this same linear adjustment is applied across all tokens, which can inadvertently distort the nuanced, relative attention patterns among the relevant tokens that are crucial for complex reasoning. In essence, the model is forced to choose between effective noise filtering and preserving its reasoning capacity.
@@
  • We reveal that the inherent triple-wise nature of the noise filtering process may necessitate numerous transformer layers for effective document preprocessing, which means we can hardly filter out noise documents with limited number of transformer layers.
  • We show that fine-tuning the model to filter out noise while effectively taking advantage of the relevant information is challenging
@@ 2 Related Work
RAG in LLM. Many recent works have explored better RAG strategies. RAG_replug treat the model as a black box and design a retriever to enhance performance. However, researchers have identified that noise in the prompt can negatively impact performance (RAG_distraction; retrieval_noise_25_1; RAG_robust_finetune_25), some researchers focus on eliminating this noisy information and compress the noisy documents(longllmlingua; long_context_llm; long_context_medical; prompt_compression). And some others try to fine-tune the large language model to make it more robust to noisy information(RAG_robust; RAG_RAFT; RAG_noise_finetune_25).
@@ 3 The Triple-Wise Problem
Clearly, noise in the documents adversely affects the performance of large language models (LLMs). And various researchers are trying to develop new retrieval methods to reduce the number of noisy information, and some others try to develop a filtering model to filter out the irrelevant information before input to LLM. In this section, we show that filtering out noise documents is inherently a complex problem, limited number of transformer layers can hardly solve the problem.
For an input sequence ${\bm{X}}=[{\bm{x}}_{0}^{T},{\bm{x}}_{1}^{T},\ldots,{\bm{x}}_{n-1}^{T}]$, the vector $\bm{r}$} indicates the relevance of each token. Specifically, for each token ${\bm{x}}_{i}$, a relevance score of $r_{i}=0$ signifies that the token is irrelevant to the query. It’s crucial to note that calculating $r_{i}$ does not solely depend on the token ${\bm{x}}_{i}$ and the query; rather, it may require the involvement of three or more tokens. For instance, in the query “Alice is exhausted, but Bob is still very excited, showing no signs of fatigue. How does Bob feel?”, the word “exhausted” acts as noise and should be excluded during inference. However, determining relevance necessitates considering “Bob” in the query alongside “Alice”, which is the subject of “exhausted”, only in this way the model can understand that “exhausted” describes another person. Therefore, identifying the relevance of a token demands information from multiple tokens, yet self-attention computes relationships only between pairs, making it challenging to address this issue within a single transformer layer.
@@
Consequently, a single layer of multi-head attention struggles to assess the relevance of a token. As noted in 3Match, To effectively solve the triple-wise problem, we need the multi-layer transformer to have width, depth, embedding dimension, or bit complexity at least $N^{\Omega(1)}c$, where $N$ is context length and $c$ is a constant represents the embedding dimension required to represent noise filtering information. This is impractical for retriever models as it only shows limited depth, width and embedding dimension.
The challenge arises from the triple-wise nature of the problem contrasted with the pairwise nature of attention; the model can only evaluate a token’s relevance when its embedding contains substantial information about the whole context. For example, as shown in Figure 1, to assess the token “exhausted” in the sentence “Alice is exhausted. How does Bob feel?”, the embedding must encompass information about its subject, “Alice/Bob”, as well as contextual details from phrases like “is/isn’t” and “feeling”. Therefore, a significant amount of information must be incorporated into the embedding before any judgment can be made, it actually requires the understanding about the meaning of the whole sentence before judging the relevance of the token.
However, a single layer of self-attention can only consider the input $\sum_{j}a_{i,j}{\bm{x}}_{j}$, which contains information up to $mp$, where $m$ represents the embedding dimension and $p$ indicates precision. The term $mp$ signifies the maximal entropy of the embedding, which means the maximal information the embedding can carry. But this can hardly represent the meaning of the whole sentence, especially when the input sequence is long.
This indicate that current retrieval methods like contriever (Contriever) would like to incorporate noisy information in the retrieved documents. However, current large language models holds great embedding dimension (4096 for Llama 3 8B), which should be able to solve the triple-wise problem theoretically. This suggests the burden of noise filtering should shift from the limited-capacity retriever to the powerful LLM itself. However, as we will show next, making the LLM robust is not a straightforward task.
@@ 4 Robustness of LLM When Faced with Noise
Clearly, noise in the documents negatively impacts the performance of LLMs, and noisy information is inevitable in RAG. Therefore some researchers focus on fine-tuning the model to make it more robust when faced with noisy information thereby enhancing performance (RAG_robust; RAG_RAFT; longllmlingua). And there is a question that, since filtering out noise is inherently a difficult problem, can the fine-tuned LLM effectively filter out irrelevant information and gather useful information to get the final answer?
Let $\bm{r}$} represents the relevance of tokens, $r_{i}=0$ means that token ${\bm{x}}_{i}$ is a noise token, otherwise the token is relevant. Let $attn({\bm{x}}_{i},{\bm{x}}_{j})=({\bm{W}}_{q}{\bm{x}}_{i})^{T}{\bm{W}}_{k}{\bm{x}}_{j}$ represent the attention pattern which is trained to extract relevant information. And we want to filter out irrelevant information as following:
@@
where $\sigma$ means the softmax function. This shows that the desired attention pattern should effectively exclude noise while preserving the attention pattern of relevant tokens. Therefore, $\widehat{attn}$ can be considered the optimal response when confronted with noise, as it effectively filters out irrelevant tokens and utilizes the relevant information in the most efficient manner.
@@
Apparently, if we want to approximate the optimal reasoning pattern, we need $\xi_{r}\approx 0$, so $\xi_{r}\lesssim\ln\frac{1}{1-\epsilon}\approx 0$. This shows that, to effectively fine-tune the model for filtering out irrelevant information in the attention matrix, we need $\xi_{r}\approx 0$. This implies that for all tokens to be retained, ${\bm{x}}_{i}^{T}\Delta{\bm{W}}{\bm{x}}_{j}$ must remain nearly constant. A solution could make $\Delta{\bm{W}}||$ small or even 0, but this actually also requires to split the tokens to be retained and to be filtered, so a small $\Delta{\bm{W}}||$ does not solve the problem as shown in the Appendix. As a result, approximating the optimal solution proves to be quite challenging because the attention will be disturbed due to the noise filtering fine-tuning.
This shows that if we want to maintain the attention pattern, then the extra attention score must be close, so it can remain consistent after softmax. Also, although the analysis mainly focus on attention module, multi head attention simply fuse the information of different heads. It does not modify the self attention mechanism so our analysis still holds, about the information fusing linear layers (MLPs), we will show that the MLP layers also can not filter out the noisy information in the following, so our analysis also satisfies in multi head attention.
Since the self-attention layer struggles to filter out irrelevant information while maintaining optimal reasoning performance, a pertinent question arises: can the feed-forward network connected to the self-attention layer effectively filter out mistakenly incorporated information? Here, we denote the information necessary to assess the relevance of the involved tokens as $\hat{{\bm{w}}}$, so $H(\hat{{\bm{w}}})=H({\bm{w}}_{i1},{\bm{w}}_{i2},\ldots)$, where ${\bm{w}}_{ij}$ represents the relevance information of the token aggregated by self-attention and those tokens form the input to the feed forward layer.
However, it is difficult to integrate additional information about $\hat{{\bm{w}}}$ (i.e., to increase $I(\hat{{\bm{w}}};{\bm{v}})$) because the input to the MLP layer is given by ${\bm{v}}=\sum_{j}a_{i,j}{\bm{x}}_{j}{\bm{W}}_{v}$, resulting in $\bm{v}$} being an embedding of dimension $m$ and precision $p$. This limited dimensionality makes it challenging to encapsulate all the necessary information for filtering out irrelevant documents for all involved tokens. Consequently, the Feed Forward Layer is unlikely to achieve optimal performance when dealing with noisy input.
Also, as show in IB_generalization, when the model is trying to filter out noise information, but some semantic information is also incorporated, then it will downweight the performance on the filtering process and vice versa. This directly means that the model can not effectively filter out noise information while taking advantage of the semantic information.
Consequently, the FFN is caught in a difficult trade-off. It cannot be an expert filter and an expert reasoner at the same time when its input is already contaminated. Relying on the FFN to clean up the mess made by the attention layer is an inefficient strategy that compromises the model’s overall performance. We show more analysis in Appendix B.2
@@ 5.1 Attention Rectification
Conventional fine-tuning paradigms, primarily adjust the model’s behavior by introducing an update matrix, $\Delta{\bm{W}}$, to the original weight matrix $\bm{W}$}. The modified attention score between a query token $i$ and a key token $j$ is computed based on their embeddings, ${\bm{x}}_{i}$ and ${\bm{x}}_{j}$. However, as noted in Theorem 4.1, a simple linear addition of the update term ${\bm{x}}_{i}^{T}\Delta{\bm{W}}{\bm{x}}_{j}$ faces a fundamental trade-off: it struggles to simultaneously (1) create a sufficiently large margin between the attention scores of relevant and irrelevant tokens, and (2) preserve the nuanced, relative attention patterns among multiple relevant tokens, which is crucial for the model’s intrinsic reasoning capabilities.
@@
  • Sharp Discrimination: The steep gradient of $tanh(x)$ around $x=0$ creates a sharp transition, effectively amplifying the margin between positive (potentially relevant) and negative (irrelevant) attention updates. This serves as a powerful mechanism for filtering out noise.
@@
From a regularization standpoint, our method mitigates overfitting to noisy signals in the training data. Token embeddings can contain spurious correlations. An unconstrained linear update might learn to assign attention with high variance scores based on such noise. The saturating nature of the tanhtanh component in $g(x)$ acts as a "soft clamp" constraining these potentially extreme values within a bounded range ($\xi$). This restriction prevents the model from becoming overly sensitive to noisy features, thereby enhancing its robustness and generalization performance. For instance, an attention update that might vary between $\xi\pm\delta$ due to noise would be regularized to $\xi\pm\epsilon$, where $\epsilon<<\delta$.
@@
5.2 The Auto-Regressive Nature Makes it Harder
Judging the relevance of token is inherently a difficult problem, small models can hardly solve the problem, therefore, we can only try to make the LLM more robust to noisy information, this is possible because LLMs hold more depth, width and embedding dimension.
However, most LLMs are trained in an auto-regressive manner, meaning that a token cannot aggregate information from any tokens that appear later in the sequence; thus, $a_{i,j}=0$ if $j>i$. Typically, the query is positioned after the document tokens, preventing these document tokens from assessing their relevance effectively because they have no access to the query. Consequently, the relevance judgment must occur during the calculation of the query embeddings, which is usually much shorter than the document, making the nose filtering harder because we need to judge the relevance of $n_{doc}$ tokens during the calculation of $n_{query}$ tokens.
Instead, if we position the query ahead of documents, then the relevance can be effectively calculated during the calculation document token embedding. This arrangement enables the information of query to be effectively transferred to the document tokens for relevance judgment. In this way we can judge the relevance of $n_{doc}$ tokens during the calculation of $n_{doc}$ tokens, as $n_{doc}$ is usually mucn larger than $n_{query}$, so placing the query ahead could help. Therefore, we can hypothesize that when there is noise in the document, placing the query at the beginning would help the judgment. We conduct experiments on various datasets in Appendix A to show that placing the query ahead can effectively help the performance.
@@ 5.2 The Auto-Regressive Nature Makes it Harder
$\xi=10$ 66.9 66.4 74.1 73.4 87.2 85.8 97.9 97.4 47.3 46.2 74.3
Table 1: Performance of our fine-tuning method when faced with explicit distracting documents. It is worth noting that $\xi=0$ results in $g(x)=x$
@@ 6.1 Datasets and Metrics
To evaluate the performance of our proposed method, we use Nature Questions (natural_question), TriviaQA (TriviaQA) which is traditionally used to evaluate the noise robustness of RAG system. Also, we use multi hop reasoning datasets HotpotQA (hotpotqa) and 2WikiMultiHopQA (2wikimultihop) as well as the long form QA dataset ASQA (ASQA) to show the performance of our method. For the first 4 datasets, we use exact match to measure the performance which is determined by whether the predicted answer contains the ground-truth answer. For ASQA, we measure the percentage of short answers are shown in the generated answer to evaluate the performance.
For all 5 datasets, we use Dense Passage Retriever as the retriever and retrieve some documents and select the first 3 documents that are not presented in the gold documents as the noisy documents, then combine the noisy documents with the gold documents as the input to the LLM, showing that our method can effectively distinguish distracting documents.
For the first 4 datasets, we randomly select 3000 samples to test and another 7000 to train. For ASQA we use the split of ALCE (ALCE) and use the 948 samples to test the performance and another 4000 for training.
More Experimental settings can be seen in the Appendix A
Table 2: The performance for Qwen2.5-7B and Mistral-7B with our activation function
NQ TriviaQA HotpotQA 2wiki ASQA
Qwen 7B vanilla 58.5 54.1 71.0 65.7 42.4
@@
ours 62.7 69.9 84.3 96.3 48.3
Mistral 7B vanilla 59.8 64.0 77.4 78.3 46.1
@@
ours 66.3 72.4 86.7 96.8 51.3

### v2

@@ 1 Introduction
  • Filtering out irrelevant information is inherently difficult, small retrieval models fail to solve it.
  • Fine-tuning the LLM can hardly distinguish irrelevant information while taking advantage of relevant ones.
About the difficulty of filtering out irrelevant information. Considering the query ‘Alice and Bob are running, Bob is exhausted. How does Bob feel?’ Here, assessing the relevance of ‘exhausted’ necessitates considering the token ‘Bob’, ‘feel’ in the query alongside ‘Bob’ which is the subject of ‘exhausted’. Thus, evaluating relevance requires the information from three or even more tokens. Yet, the attention mechanism typically computes only pairwise relationships, making it challenging to resolve this issue within a limited number of transformer layers (3Match). Therefore filtering out noise documents is inherently difficult which explains why current small retriever models (Contriever; BM25; DPR) would always incorporate some noisy information in the retrieval results.
Therefore, a question is can powerful LLMs solve this problem. We hope that the LLM could be robust to noisy information while extracting helpful information (RAG_robust; RAG_robust_finetune_25). However, we argue that standard fine-tuning is structurally ill-suited for this task. The core issue lies in a fundamental trade-off imposed by their linear update mechanism. To filter noise, the learned parameter update must apply a strong negative adjustment to the attention scores of irrelevant tokens. Yet, this same linear adjustment is applied across all tokens, which can inadvertently distort the nuanced, relative attention patterns among the relevant tokens that are crucial for complex reasoning. In essence, the model is forced to choose between effective noise filtering and preserving its reasoning capacity.
@@
  • We reveal that the inherent triple-wise nature of the noise filtering process may necessitate numerous transformer layers, which means we can hardly filter out noise documents with small models.
  • We show that fine-tuning the LLM to filter out noise while effectively taking advantage of the relevant information is challenging
@@ 2 Related Work
RAG in LLM. Many recent works have explored better RAG strategies. RAG_replug treat the model as a black box and design a retriever to enhance performance. However, researchers have identified that noise in the context can negatively impact performance (RAG_distraction; retrieval_noise_25_1; RAG_robust_finetune_25), some researchers focus on eliminating this noisy information and compress the noisy documents(longllmlingua; long_context_llm; long_context_medical; prompt_compression). And some others try to fine-tune the large language model to make it more robust to noisy information(RAG_robust; RAG_RAFT; RAG_noise_finetune_25).
@@ 3 The Triple-Wise Problem
Clearly, noise in the documents adversely affects the performance of large language models (LLMs). And various researchers are trying to develop new retrieval methods to reduce the number of noisy information, and some others try to develop a filtering model to filter out the irrelevant information before input to LLM. In this section, we show that filtering out noise documents is inherently a complex problem. It is difficult to filter out noise by small model based rerankers/retrievers.
For input ${\bm{X}}=[{\bm{x}}_{0}^{T},{\bm{x}}_{1}^{T},\ldots,{\bm{x}}_{n-1}^{T}]$ with the query and document, and we should decide is the document relevant to query. It’s crucial to note that calculating the relevance of the document require the involvement of many tokens. For instance, in the query “Alice and Bob are running, Bob is exhausted. How does Bob feel?”, the context “exhausted” describes the feeling of Bob, which serves as a useful contenxt. However, determining relevance necessitates considering “Bob”, “feel” in the query alongside “Bob”, which is the subject of “exhausted”, only in this way the model can understand that “exhausted” describes Bob and the query asks for the feeling of Bob. We show it more clearly in Figure 1. Therefore, identifying the relevance of a token demands information from multiple tokens. However, self-attention computes relationships only between pairs, making it challenging to effectively address this issue.
@@
As self attention only calculates pair-wise relationship between tokens, and judging the relevance of the document requires the involvement of multiple tokens. It is necessary to stack multiple attention layers to aggregate information from different tokens to conduct judgment. However, as noted in 3Match, To effectively solve the triple-wise problem, we need the multi-layer transformer to have width, depth, embedding dimension, or bit complexity at least $N^{\Omega(1)}c$, where $N$ is context length and $c$ is a constant represents the embedding dimension required to represent noise filtering information. This is impractical for small models as it only shows limited depth, width and embedding dimension.
The challenge arises from the triple-wise nature of the problem contrasted with the pairwise nature of attention; the model can only evaluate a token’s relevance when its embedding contains substantial information about the whole context. For example, as shown in Figure 1, to assess the token “exhausted” in the sentence “Bob is exhausted. How does Bob feel?”, the embedding must encompass information about its subject of “exhausted” and the subject of “feel”, “Bob”, as well as contextual details from phrases like “is/isn’t” and “feeling”. Therefore, a significant amount of information must be incorporated into the embedding before any judgment can be made, it actually requires the understanding about the meaning of the whole sentence before judging the relevance of the token.
However, a single layer of self-attention can only consider the input of fixed dimension embedding, which contains information up to $mp$, where $m$ represents the embedding dimension and $p$ indicates precision of embedding. The term $mp$ signifies the maximal information carried by the embedding. But this can hardly represent the meaning of the whole sentence, especially when the input sequence is long.
This indicates that trying to filter out irrelevant context is difficult, the input to LLM would like to incorporate noisy information in the retrieved documents. However, current large language models holds great embedding dimension (4096 for Llama3-8B) and depth (32 for Llama3-8B), which should be able to solve the triple-wise problem theoretically. This suggests the burden of noise filtering should shift from the limited-capacity retriever to the powerful LLM itself. However, as we will show next, making the LLM robust is not a straightforward task.
@@ 4 Robustness of LLM When Faced with Noise
Clearly, noise in the documents negatively impacts the performance of LLMs, and noisy information is inevitable in RAG. Therefore some researchers focus on fine-tuning the model to make it more robust when faced with noisy information thereby enhancing performance (RAG_robust; RAG_noise_finetune_25; RAG_robust_finetune_25; longllmlingua). And there is a question that, can the fine-tuned LLM effectively filter out irrelevant information and gather useful information to get the final answer?
Let $\bm{r}$} represents the relevance of tokens, $r_{i}=0$ means that token ${\bm{x}}_{i}$ is a noise token, otherwise the token is relevant. Let $attn({\bm{x}}_{i},{\bm{x}}_{j})=({\bm{W}}_{q}{\bm{x}}_{i})^{T}{\bm{W}}_{k}{\bm{x}}_{j}$ represent the attention pattern which is trained to extract relevant information.
And we want to filter out irrelevant information while preserving the attention pattern on relevant tokens, this can be represented as following:
@@
where $\sigma$ means the softmax function. This shows that the desired attention pattern should effectively exclude noise while preserving the attention pattern of relevant tokens.
Therefore, $\widehat{attn}$ can be considered the optimal response when confronted with noise, as it effectively filters out irrelevant tokens and utilizes the relevant information in the most efficient manner.
@@
Apparently, if we want to approximate the optimal reasoning pattern, we need $\xi_{r}\approx 0$, so $\xi_{r}\lesssim\ln\frac{1}{1-\epsilon}\approx 0$. This shows that, to effectively fine-tune the model to filter out irrelevant information in the attention matrix, we need $\xi_{r}\approx 0$. This implies that for all tokens to be retained, ${\bm{x}}_{i}^{T}\Delta{\bm{W}}{\bm{x}}_{j}$ must remain nearly constant. A solution could make $\Delta{\bm{W}}||$ small or even 0, but this actually also requires to split the tokens to be retained and to be filtered, so a small $\Delta{\bm{W}}||$ does not solve the problem as shown in Appendix B.2. As a result, approximating the optimal solution proves to be quite challenging because the attention will be disturbed due to the noise filtering fine-tuning.
One might argue that preserving the original attention pattern is unnecessary. Perhaps the model could learn a new, superior attention pattern during fine-tuning that excels at both filtering noise and leveraging semantic information. However, this perspective overlooks a fundamental conflict between these two objectives.
To conduct the two task, we require the input to attention contain two different kinds of information, noise filtering and semantic reasoning. As suggested by IB_generalization, a model with finite capacity attempts to optimize for two different tasks would show suboptimal performance on both task because the two different tasks requires different information, and each act as noise to each other, so when the model is trying to filter out noise, but some semantic information is also incorporated into the input, then it will downweight the performance on the filtering process and vice versa.
This also fits for the feed forward layers. It cannot be an expert filter and an expert reasoner at the same time when its input is already contaminated. Therefore, relying on the FFN to clean up the mess made by the attention layer is an inefficient strategy that compromises the model’s overall performance. We show more analysis in Appendix B.3
It is worth noting that although the analysis mainly focus on single head attention, for multi head attention, we can use one head focus on filtering out noise and another head focus on taking advantage of the relevant information. This is indeed ideal, but fine-tuning based on existing LLMs fails to greatly influence the existing parameters, otherwise it will cause catastrophic forgetting (catastrophic_forgetting1; catastrophic_forgetting2).
@@ 5.1 Attention Rectification
Conventional fine-tuning paradigms, primarily adjust the model’s behavior by introducing an update matrix, $\Delta{\bm{W}}$, to the original weight matrix $\bm{W}$}. The modified attention score between a query token $i$ and a key token $j$ is computed based on their embeddings, ${\bm{x}}_{i}$ and ${\bm{x}}_{j}$. However, as noted in Theorem 4.1, a simple linear addition of the update term ${\bm{x}}_{i}^{T}\Delta{\bm{W}}{\bm{x}}_{j}$ faces a fundamental trade-off: it struggles to simultaneously (1) create a sufficiently large margin between the attention scores of relevant and irrelevant tokens, and (2) preserve and optimize the nuanced, relative attention patterns among multiple relevant tokens, which is crucial for the model’s intrinsic reasoning capabilities.
@@
  • Sharp Discrimination: The steep gradient of $tanh(x)$ around $x=0$ creates a sharp transition, effectively amplifying the margin between positive (potentially relevant) and negative (irrelevant) attention updates. This serves as a powerful mechanism for filtering out noise as a large negative attention could be allocated to unrelated tokens.
@@
From a regularization standpoint. We actually require the attention module to conduct two tasks, nose filtering and relevant information aggregation. The two tasks may require different information to process and they are actually noise to each other. In this way, an unconstrained linear update might learn to assign attention with high variance scores based on such noise as shown in IB_generalization. The saturating nature of the tanhtanh component in $g(x)$ acts as a "soft clamp" constraining these potentially high variance values within a bounded range ($\xi$). This restriction prevents the model from becoming overly sensitive to noisy features, thereby enhancing its robustness and generalization performance. For instance, an attention update that might vary between $\xi\pm\delta$ with only linear update, and with the tanh activation, it would be regularized to $\xi\pm\epsilon$, where $\epsilon<<\delta$.
@@
5.2 The Auto-Regressive Nature may Cause Problem
With the activation, the attention module can effectively filter out noise information. However, most LLMs are trained in an auto-regressive manner, meaning that a token cannot aggregate information from any tokens that appear later in the sequence; thus, $a_{i,j}=0$ if $j>i$. Typically, the query is positioned after the document tokens, preventing these document tokens from assessing their relevance effectively because they have no access to the query. Consequently, the relevance judgment must occur during the calculation of the query embeddings, which is usually much shorter than the document, making the nose filtering harder because we need to judge the relevance of $n_{doc}$ tokens during the calculation of $n_{query}$ tokens.
Instead, if we position the query ahead of documents, then the relevance can be effectively calculated during the calculation document token embedding. This arrangement enables the information of query to be effectively transferred to the document tokens for relevance judgment. In this way we can judge the relevance of $n_{doc}$ tokens during the calculation of $n_{doc}$ tokens, as $n_{doc}$ is usually much larger than $n_{query}$, so placing the query ahead could help. Therefore, we can hypothesize that when there is noise in the document, placing the query at the beginning would help the judgment. We conduct experiments on various datasets in Appendix A to show that placing the query ahead can effectively help the performance.
@@ 5.2 The Auto-Regressive Nature may Cause Problem
$\xi=10$ 66.9 65.4 74.1 73.4 87.2 85.8 97.9 97.4 47.3 46.2 74.2
Table 1: Performance of our fine-tuning method when faced with explicit distracting documents. reverse means we place the query ahead of documents and vanilla means the query is placed after documents
@@ 6.1 Datasets and Metrics
To evaluate the performance of our proposed method, we use Nature Questions (natural_question), TriviaQA (TriviaQA) which is traditionally used to evaluate the noise robustness of RAG system. Also, we use multi hop reasoning datasets HotpotQA (hotpotqa) and 2WikiMultiHopQA (2wikimultihop) as well as the long form QA dataset ASQA (ASQA) to show the performance of our method. For the first 4 datasets, we use accuracy to measure the performance which is determined by whether the predicted answer contains the ground-truth answer. For ASQA, we measure the percentage of short answers are shown in the generated answer to evaluate the performance.
For all 5 datasets, we use Dense Passage Retriever (DPR) as the retriever, we retrieve some documents and select the first 3 documents that are not presented in the gold documents as the noisy documents, then combine the noisy documents with the gold documents as the input to the LLM, showing that our method can effectively distinguish distracting documents while taking advantage of relevant ones.
For the first 4 datasets, we randomly select 3000 samples to test and another 7000 to train. For ASQA we use the split of ALCE (ALCE) and use the 948 samples to test the performance and another 4000 for training. More Experimental settings can be seen in Appendix A
NQ Trivia Hotpot 2wiki ASQA
Qwen 7B vanilla 53.5 59.9 74.8 67.1 43.5
@@
tanh 62.7 69.9 84.3 96.3 48.3
Mistral 7B vanilla 52.3 63.5 71.6 58.6 46.0
@@
tanh 66.3 72.4 86.7 96.8 51.3
Table 2: The performance for Qwen2.5-7B ($\xi=5$) and Mistral-7B ($\xi=3$) with our activation function